OpenStudy (anonymous):

You place a cup of 200 degrees F coffee on a table in a room that is 67 degrees F, and 10 minutes later, it is 195 degrees F. Approximately how long will it be before the coffee is 180 degrees F? Use Newton's law of cooling: T(t)=T[a]+(T[o]-T[a])e^-kt A. 40 minutes B. 15 minutes C. 1 hour D. 35 minutes

OpenStudy (anonymous):

PLEASE HELP !! i cant get this wrong .. i really need to know the process

OpenStudy (anonymous):

and answer ...

OpenStudy (anonymous):

I had a similar question to this... in order to find the t, you need to find the a and the k values

OpenStudy (anonymous):

mmmhhhh ok

OpenStudy (anonymous):

T(0) = a + 67 = 190, solve for a

OpenStudy (anonymous):

123

OpenStudy (anonymous):

T(30) = 123e^10k + 70 = 195, now solve for k

OpenStudy (anonymous):

ok is it ok when im done solving it can u check if i got the answer right?? @jtryon

OpenStudy (anonymous):

yes, I will check for you

OpenStudy (anonymous):

thank you n_n @jtryon

OpenStudy (anonymous):

wait, I made an error

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

you want to use T(30) = 123e^10k + 67 = 195, because 67 is the surrounding air temp

OpenStudy (anonymous):

from the problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait im confuse do i subtract 67 from 195 ?? cuz the T(30) is confusing me @jtryon

OpenStudy (anonymous):

yes, so you would solve it like any other problem and subtract 67 from both sides

OpenStudy (anonymous):

oh okk soo just ignore the T(30)

OpenStudy (anonymous):

yes, that is just given information in the problem

zepdrix (zepdrix):

The initial temperature of the coffee \(\Large\rm T_o\) is 200. The ambient temperature \(\Large\rm T_a\) is 67. (I think that's what the a stands for at least..). And they tell us at time \(\Large\rm t=10\), the temperature has dropped to 195. \(\Large\rm T(10)=195\). \[\Large\rm T(t)=T_a+(T_o-T_a)e^{-kt}\]So you can plug some stuff in and solve for k, yah? Hmm I can't figure out where this T(30) is coming from that you guys are using :(

OpenStudy (anonymous):

@zepdrix, I just used that as a placeholder for the equation that I want to solve

OpenStudy (anonymous):

but what you are solving is 123e^10k + 67 = 195, for k

zepdrix (zepdrix):

Oh :)

OpenStudy (anonymous):

ohhh ok then .... getting confusedd

OpenStudy (anonymous):

go back to the original equation, you want to find k so that you can plug that in to find t

OpenStudy (anonymous):

have you figured out what k is?

zepdrix (zepdrix):

Woops* \[\Large\rm 1\color{red}{3}3e^{\color{red}{-}10k} + 67 = 195\]^I think this is the equation, yes? Hope I did my math correctly >.<

OpenStudy (anonymous):

I got 123e

OpenStudy (anonymous):

\[123e ^{10k} = 128 \]

OpenStudy (anonymous):

thats where i am at ...

OpenStudy (anonymous):

It's not going to be 128, less than that

OpenStudy (anonymous):

so we know that a = 123

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your equation would be 123e^10k + 67 (Ts) = 195 (T0)

OpenStudy (anonymous):

hoping you are not confused by T sub-zero and t sub-s

OpenStudy (anonymous):

if we subtract 67 from both sides, we get 128 (you are right up to here)

OpenStudy (anonymous):

not yet hahaha

OpenStudy (anonymous):

I forgot to bring my 1 down

zepdrix (zepdrix):

\[\Large\rm T(t)=T_a+(T_o-T_a)e^{-kt}\]\[\Large\rm T(t)=67+(200-67)e^{-kt}\]\[\Large\rm T(t)=67+133e^{-kt}\] I'm not sure where you got 123 :P So weird...

OpenStudy (anonymous):

what would that equation lead me to ??? @zepdrix

OpenStudy (anonymous):

Oh, it is 133 :P

OpenStudy (anonymous):

so once you have the a value of 133, you plug that into this equation 133e^10k+67 = 195

OpenStudy (anonymous):

and you are solving for the k value

OpenStudy (anonymous):

do you know the steps to solve that equation?

OpenStudy (anonymous):

\[133e ^{10k} = 128\]

OpenStudy (anonymous):

thats as far as i go ...

OpenStudy (anonymous):

that is right so far

zepdrix (zepdrix):

minus in front of the k, yes? :o

OpenStudy (anonymous):

hmm, I don't have a minus in front of the k

OpenStudy (anonymous):

the formula is Ae^kt + Ts

OpenStudy (anonymous):

@zepdrix, where do you get a minus k?

zepdrix (zepdrix):

From the formula he posted :o

OpenStudy (anonymous):

what did you get for k, @meln_n

OpenStudy (anonymous):

Oh I'm using the formula T= T(t) = ae^kt + Ts

OpenStudy (anonymous):

i cant solve it im still stuck at \[133e ^{10k} = 195 \]

zepdrix (zepdrix):

Since the coffee temperature is `higher` than the ambient temperature, I think it makes sense to have a negative k, as time increases, our temperature falls, getting closer to the ambient temperature.

OpenStudy (anonymous):

So you are going to have 133e^-10k = 128

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@zepdrix, that look right?

OpenStudy (anonymous):

using the formula that he posted

zepdrix (zepdrix):

Yah seems like it :D Solve for k! You get a really ugly decimal.

zepdrix (zepdrix):

Understand how to isolate the k?

OpenStudy (anonymous):

You have to divide by 133 on both sides

OpenStudy (anonymous):

no i dont thats where i am stuck ... am i suppose to divide next @zepdrix

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

Do you know what to do after that?

OpenStudy (anonymous):

\[e ^{-10k} = .9624\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Take the natural log of both sides

OpenStudy (anonymous):

soooo \[\log e ^{-10k} = \log.9624 \]

OpenStudy (anonymous):

would it turn to ln

OpenStudy (anonymous):

yes, and then what decimal do you get?

OpenStudy (anonymous):

mmmmhhh i got something weird -.038325

OpenStudy (anonymous):

what did you plug in?

OpenStudy (anonymous):

wait i got .00383

OpenStudy (anonymous):

is it still wrong

zepdrix (zepdrix):

You're on the right track. Then you need to divide both sides by -10 to solve for k.

OpenStudy (anonymous):

\[\log e ^{-10k} = \log.9624 \]

OpenStudy (anonymous):

from that ^^^ i divide -10 ??? how soo

OpenStudy (anonymous):

so you have log .9624 / -10

OpenStudy (anonymous):

yes i have that !

OpenStudy (anonymous):

\[\frac{ \ln.9624 }{ -10 }\]

OpenStudy (anonymous):

@zepdrix, I think he has the nasty decimal :P

OpenStudy (anonymous):

but when you divide by a negative, you get a positive decimal

zepdrix (zepdrix):

Oh sorry, forgot to explain I guess :) \[\Large\rm \ln(e^{-10k})\]Apply rule of exponents, bring the -10k in front,\[\Large\rm -10k \ln(e)\]Then the ln(e) is simply 1,\[\Large\rm -10k(1)=-10k\]

OpenStudy (anonymous):

so what did you get for your k?

OpenStudy (anonymous):

i got .00383

zepdrix (zepdrix):

So plug your k into your formula:\[\Large\rm T(t)=67+133e^{-\color{orangered}{k}t}\](Or wait til the very end of the problem to plug it in, that might be easier). They want to know what time value gives us a temperature of 180.\[\Large\rm 180=67+133e^{-\color{orangered}{k}t}\]Solve for t

OpenStudy (anonymous):

You have to do some more steps in order to find the time value

OpenStudy (anonymous):

Do you know how to get t alone?

OpenStudy (anonymous):

not exactly i did something and i got to here

OpenStudy (anonymous):

\[e ^{-kt} = .8496\]

OpenStudy (anonymous):

So what I have for the equation is 180 = 67 + 133e^-.00383t

OpenStudy (anonymous):

ok i used that i just havent input the k value that we found

OpenStudy (anonymous):

@zepdrix, so I have 113 = 133e^-.00383t

OpenStudy (anonymous):

would it be like this then \[e ^{-.00383t} = .8496\]

OpenStudy (anonymous):

what happened to the rest of the equation?